# Dense $\mathbb{X}$ Retrieval: What Retrieval Granularity Should We Use?

### v2

@@ Appendix E Additional Results
Decompose the "Content" into clear and simple propositions, ensuring they are interpretable out of context. 1. Split compound sentence into simple sentences. Maintain the original phrasing from the input whenever possible. 2. For any named entity that is accompanied by additional descriptive information, separate this information into its own distinct proposition. 3. Decontextualize the proposition by adding necessary modifier to nouns or entire sentences and replacing pronouns (e.g., "it", "he", "she", "they", "this", "that") with the full name of the entities they refer to. 4. Present the results as a list of strings, formatted in JSON. Input: Title: Ēostre. Section: Theories and interpretations, Connection to Easter Hares. Content: The earliest evidence for the Easter Hare (Osterhase) was recorded in south-west Germany in 1678 by the professor of medicine Georg Franck von Franckenau, but it remained unknown in other parts of Germany until the 18th century. Scholar Richard Sermon writes that "hares were frequently seen in gardens in spring, and thus may have served as a convenient explanation for the origin of the colored eggs hidden there for children. Alternatively, there is a European tradition that hares laid eggs, since a hare’s scratch or form and a lapwing’s nest look very similar, and both occur on grassland and are first seen in the spring. In the nineteenth century the influence of Easter cards, toys, and books was to make the Easter Hare/Rabbit popular throughout Europe. German immigrants then exported the custom to Britain and America where it evolved into the Easter Bunny." Output: [ "The earliest evidence for the Easter Hare was recorded in south-west Germany in 1678 by Georg Franck von Franckenau.", "Georg Franck von Franckenau was a professor of medicine.", "The evidence for the Easter Hare remained unknown in other parts of Germany until the 18th century.", "Richard Sermon was a scholar.", "Richard Sermon writes a hypothesis about the possible explanation for the connection between hares and the tradition during Easter", "Hares were frequently seen in gardens in spring.", "Hares may have served as a convenient explanation for the origin of the colored eggs hidden in gardens for children.", "There is a European tradition that hares laid eggs.", "A hare’s scratch or form and a lapwing’s nest look very similar.", "Both hares and lapwing’s nests occur on grassland and are first seen in the spring.", "In the nineteenth century the influence of Easter cards, toys, and books was to make the Easter Hare/Rabbit popular throughout Europe.", "German immigrants exported the custom of the Easter Hare/Rabbit to Britain and America.", "The custom of the Easter Hare/Rabbit evolved into the Easter Bunny in Britain and America." ] Input: <a new passage> Output:

### v3

@@ Appendix G Error Case Study
Passage $\Rightarrow$ Propositions Decompose the "Content" into clear and simple propositions, ensuring they are interpretable out of context. 1. Split compound sentence into simple sentences. Maintain the original phrasing from the input whenever possible. 2. For any named entity that is accompanied by additional descriptive information, separate this information into its own distinct proposition. 3. Decontextualize the proposition by adding necessary modifier to nouns or entire sentences and replacing pronouns (e.g., "it", "he", "she", "they", "this", "that") with the full name of the entities they refer to. 4. Present the results as a list of strings, formatted in JSON. Input: Title: Ēostre. Section: Theories and interpretations, Connection to Easter Hares. Content: The earliest evidence for the Easter Hare (Osterhase) was recorded in south-west Germany in 1678 by the professor of medicine Georg Franck von Franckenau, but it remained unknown in other parts of Germany until the 18th century. Scholar Richard Sermon writes that "hares were frequently seen in gardens in spring, and thus may have served as a convenient explanation for the origin of the colored eggs hidden there for children. Alternatively, there is a European tradition that hares laid eggs, since a hare’s scratch or form and a lapwing’s nest look very similar, and both occur on grassland and are first seen in the spring. In the nineteenth century the influence of Easter cards, toys, and books was to make the Easter Hare/Rabbit popular throughout Europe. German immigrants then exported the custom to Britain and America where it evolved into the Easter Bunny." Output: [ "The earliest evidence for the Easter Hare was recorded in south-west Germany in 1678 by Georg Franck von Franckenau.", "Georg Franck von Franckenau was a professor of medicine.", "The evidence for the Easter Hare remained unknown in other parts of Germany until the 18th century.", "Richard Sermon was a scholar.", "Richard Sermon writes a hypothesis about the possible explanation for the connection between hares and the tradition during Easter", "Hares were frequently seen in gardens in spring.", "Hares may have served as a convenient explanation for the origin of the colored eggs hidden in gardens for children.", "There is a European tradition that hares laid eggs.", "A hare’s scratch or form and a lapwing’s nest look very similar.", "Both hares and lapwing’s nests occur on grassland and are first seen in the spring.", "In the nineteenth century the influence of Easter cards, toys, and books was to make the Easter Hare/Rabbit popular throughout Europe.", "German immigrants exported the custom of the Easter Hare/Rabbit to Britain and America.", "The custom of the Easter Hare/Rabbit evolved into the Easter Bunny in Britain and America." ] Input: <a new passage> Output:
@@
Open-domain QA for LLaMA-2-7B … [demonstrations] … Refer to the passages below and answer the following question with just a few words. Title: 1972 in spaceflight. Passage: In 1972, humanity’s last crewed mission to the Moon of the 20th century was Apollo 17. Title: 1970s. Passage: Apollo 17 Astronaut Gene Cernan becomes the last man on the Moon on December 13, 1972. Title: List of Apollo missions Refer to the context above and answer the following question with just a few words. Question: when was the last time anyone was on the moon The answer is
Figure 9: Prompt for retrieval-augmented generation of open-domain QA for the LLaMA-2-7B model.